You just inherited $10,000. While you plan to squander some of it away, how much should you deposit in an account earning 7% interest per year if you'd like to have $10,000 in the account in 10 years?

Youjat heted S10 000 whie you plan to iqua der so me of airy how much should you deposit n an account eam ng 7% interest per year ส youd like to have S 10,000 in the
Answer #1

Present Worth = F(P/F, i, n)

                        = 10,000(P/F, 7%, 10)

                        = 10,000(0.5083)

                        = $5,083

Thus, you should deposit $5,083 in an account